OpenStudy (anonymous):

What characteristic describes late nineteenth-century sculptor Auguste Rodin's work?

OpenStudy (anonymous):

Rodin's work depicts figures with little expression that convey the artist's attitude toward industrialized society. Rodin's work captures a sense of movement and emotion through surface texture variations. Rodin's work features figures placed on high pedestals to distance them from the viewers. Rodin's work conveys the artist's dislike of Renaissance sculpture by Donatello and Michelangelo.

OpenStudy (anonymous):

@SoullessEyes

OpenStudy (hysusonic):

round 7 i believe

OpenStudy (anonymous):

Haha

OpenStudy (hysusonic):

so which ones can you cross out?

OpenStudy (anonymous):

A C AND D

OpenStudy (hysusonic):

there's 2 A's and an N and 2 D's? wow such logic much letters wow

OpenStudy (anonymous):

Oh no you didn't! HOLD UP WAIT A MINUTE BACK IT UP AND PUT IT IN REVERSE!!!!

OpenStudy (anonymous):

I am very logical...1+1=3 I bet you didnt know that. You may not know but i'm only 2.

OpenStudy (hysusonic):

excuse you, 1+1= window thank you very much

OpenStudy (anonymous):

oh oh....1*1=sky

OpenStudy (hysusonic):